outside the United States? Poisson distribution with Xt equal to 3: b. P(r > 3) d. Find the smallest x so that P(x s x) > 0.50. 5-56. Determine the following values associated with a
Answer #1

We are given the distribution here as:

\LARGE X \sim Poisson(3)

\LARGE P(X = 0) = e^{-3} = 0.0498

\LARGE P(X = 1) = 3e^{-3} = 0.1494

\LARGE P(X = 2) = \frac{3^2}{2}e^{-3} = 0.2240

\LARGE P(X = 3) = \frac{3^3}{3!}e^{-3} = 0.2240

\LARGE P(X = 4) = \frac{3^4}{4!}e^{-3} = 0.1680

\LARGE P(X = 5) = \frac{3^5}{5!}e^{-3} = 0.1008

a) The probability here is computed as:

\LARGE P(X \leq 3) = P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3)

\LARGE P(X \leq 3) = 0.0498 + 0.1494 + 0.2240 +0.2240

\LARGE P(X \leq 3) = 0.6472

Therefore 0.6472 is the required probability here.

b) The required probability here is computed as:

\LARGE P(X > 3) = 1 - P(X \leq 3)

\LARGE P(X > 3) = 1 - 0.6472 = 0.3528

Therefore 0.3528 is the required probability here.

c) The required probability here is computed as:

\LARGE P(2 < X \leq 5) = P(X =3 ) + P(X = 4) + P(X = 5)

\LARGE P(2 < X \leq 5) = 0.2240 + 0.1680 + 0.1008

\LARGE P(2 < X \leq 5) = 0.4928

Therefore 0.4928 is the required probability here.

d) We already computed in first part that:

\LARGE P(X \leq 3) = 0.6472

\LARGE P(X \leq 2) = 0.6472 - 0.2240 = 0.4232

Therefore the smallest value of x required here is 0.4232